what if i forgot the value of 8 times 7 how cold i use the distributive propertyto find 8 times seven
Answer by Fombitz(13823)

Let's review the distributive property.
A*(B+C) = A*B + A*C
So how does that apply to 8*7?
8*(7)=8*(6+1)=8*(5+2)=8*(4+3)
So all of these sums should be the same. Let's check. We'll work backwards.
8*(4+3) = 8*4 + 8*3 = 32 + 24 = 56
8*(5+2) = 8*5 + 8*2 = 40 + 16 = 56
8*(6+1) = 8*6 + 8*1 = 48 + 8 = 56
So then 8*7 = 56
Eventually you could get down to
8*7 = 8*1 + 8*1 + 8*1 + 8*1 + 8*1 + 8*1 + 8*1 + 8*1 or 8 added 7 times or (8 times 7) and that would still equal 56.
Multiplication is just quicker addition.